# A New Formulation for Zeroth-Order Optimization of Adversarial EXEmples in Malware Detection

### v1

@@ 3.3 Theoretical Results
In this section, we present theoretical results for ZEXE for non-convex non-smooth functions and compare them with those obtained by state-of-the-art methods. Proofs are provided in Appendix B. As we observe from Algorithm 1, given a threshold $\omega\geq 0$, for every time-step $k\in\mathbb{N}$, if $\|g_{(G_{k},h_{k})(v_{k})}\|^{2}>\omega$, then the update of the iteration is computed as in structured finite-difference methods. Moreover, as in [31], for $h>0$, $\mathbb{E}_{G}[g_{(G,h)}(\cdot)]=\nabla F_{h}(\cdot)$, where $g_{(G_{k},h_{k})}(\cdot)$ is defined in Eq. 7 and $F_{h}(v):=\mathbb{E}_{u\sim\mathbb{B}^{d}}[F_{z}(v+hu)]$ see Lemma 3 in Appendix A. Thus, for every $0\leq\underline{K}<\bar{K}$ such that for $k=\underline{K},\ldots,\bar{K}$ the condition $\|g_{(G_{k},h_{k})(v_{k})}\|^{2}>\omega$ holds, we can provide rates for the expectation of the norm of the gradient of the smooth approximation of the objective function $F_{h}$. In the following, we refer to complexity as the number of function evaluations required to achieve an accuracy $\varepsilon\in(0,1)$. We use the following notation,
$\eta_{\underline{K}}^{\bar{K}}:=\Bigg{(}\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(x_{i})\|^{2}]\Bigg{)}/A_{\underline{K}}^{\bar{K}}\qquad\text{with}\qquad A_{\underline{K}}^{\bar{K}}:=\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}.$
In particular, we perform the analysis of ZEXE for functions that satisfy the following assumption
@@
Now, we state the main theorem for the non-convex non-smooth setting.
@@
$\eta_{\underline{K}}^{\bar{K}}\leq S_{\underline{K}}^{\bar{K}}/A_{\underline{K}}^{\bar{K}}\qquad\text{with}\qquad S_{\underline{K}}^{\bar{K}}:=F_{h_{\underline{K}}}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h_{i}},$
@@
Under the same Assumptions of Theorem 3.1, let $\gamma_{k}=\gamma$ and $h_{k}=h$ with $\gamma,h>0$. Then, for $k=\underline{K},\cdots,\bar{K}$
$\eta_{\underline{K}}^{\bar{K}}\leq\frac{F_{h}(v_{\underline{K}})-\min F_{z}}{\gamma(\bar{K}-\underline{K})}+\frac{L^{3}d^{2}\sqrt{d}}{\ell h}\gamma$
Moreover, let $K=\bar{K}-\underline{K}$ and $\varepsilon\in(0,1)$. Then, if $K\geq 4(F_{h}(v_{\underline{K}})-\min F_{z})(L^{3}d^{2}\sqrt{d})\epsilon^{-2}/(\ell h)$, choosing $\gamma=\sqrt{\frac{F_{h}(v_{\underline{K}})-\min F_{z})\ell h}{KL^{3}d^{2}\sqrt{d}}}$, we have that $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$. Thus, the complexity of obtaining $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$ is of the order $\mathcal{O}(d^{2}\sqrt{d}h^{-1}\epsilon^{-2})$.
Discussion. In Theorem 3.1, we provide a rate on the expected norm of the smoothed gradient. The resulting bound consists of two parts. The first part is due to the (smoothed) functional value at the first point of the sequence for which the norm of the gradient approximation is larger than $\omega$ (i.e., the initialization). The second part is the approximation error. Recall that Assumption 1 holds, and therefore we have that $F_{h_{\underline{K}}}(v)\leq F_{z}(v)+Lh_{\underline{K}}$ for every $v\in\mathbb{R}^{d}$ see Proposition 1. Thus, by choosing the sequence $h_{k}$ small, we can reduce the gap between $F_{h}$ and $F$. However, making this sequence too small would increase the approximation error significantly. Note that with $\omega=0$, Theorem 3.1 holds for every $k\in\mathbb{N}$. In particular, we obtain a rate that differs from [31, Theorem 2] only in the dependence on the dimension in the second term of $S_{\underline{K}}^{\bar{K}}$. We can achieve the same dependence on the dimension by choosing, for instance, the sequence $h_{k}=d\bar{h}_{k}$, where $\bar{h}_{k}$ is some non-negative sequence. Moreover, note that our result is more general than [31, Theorem 2] since we do not require $h_{k}$ to be constant. In Corollary 1, we consider a constant choice for the step size and the smoothing parameter. In particular, we observe that if the number of consecutive iterations $k=\underline{K},\ldots,\bar{K}$ for which $\|g_{(G_{k},h_{k})}(v_{k})\|^{2}>\omega$ is sufficiently large, we get that $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$. This means that, under such a condition, in expectation, the iterations allow us to approach a stationary point of $F_{h}$.
@@ 4 Experiments
We now test the efficacy of ZEXE by comparing it with GAMMA, a state-of-the-art attack that injects content harvested from samples benignware [12].
@@
Baseline models. We consider two state-of-the-art machine learning Windows malware detectors. The first one is MalConv [29], an end-to-end deep convolutional neural network. MalConv takes in input programs as whole strings of bytes, cropping them if they exceed 2MB of input or, on the contrary, padding them with special value 256. Due to the absence of any notion of distance in the domain of bytes, these are processed through an 8-dimensional embedding layer that makes them comparable. The second model is GBDT-EMBER [1], which is a gradient-boosting decision tree trained a complex feature representation introduced to build the open-source EMBER dataset. In particular, EMBER features aggregates information regarding headers, imported APIs, byte histrograms, readable strings and more into 2381 features. For this paper, we leverage an open-source implementation of MalConv, with an input size shrunk to 1MB and detection threshold set to $0.5$, while for GBDT-EMBER we use its original open-source implementation, [^3] https://github.com/elastic/ember, with detection threshold set at $0.8$. Both models are wrapped to be attacked through the SecML-Malware Python library [11].
@@
Setup of ZEXE. We set the number of explored directions $\ell=5$ and the threshold on the gradient norm $\omega=10^{-4}$. We set a number of queries of $1000$, thus the number of iterations of ZEXE is $T=1000/(\ell+1)$. As practical manipulation, we use section injection, which includes new content inside input Windows programs, like already-initialized strings, resources, or other assets [12]. We initialize the manipulation with $d$ random printable bytes, ranging from the white-space character (encoded as 32) to the tilde character "~" (encoded as 126). We define the encoding mapping $\psi$ as the function that maps these bytes to their corresponding values in $\mathbb{R}^{d}$. The decoding function $\xi$ projects the perturbed manipulation onto the hypercube $[32,126]^{d}\subseteq\mathbb{R}^{d}$. It then takes the integer part of each element in the perturbed manipulation and maps those integers back to the corresponding bytes, ensuring that the bytes remain within the range of printable characters. Finally, the manipulation function $\rho$ modifies the newly injected sections. We defined the sampling space $V$ as the hypercube $[32,126]^{d}\subseteq\mathbb{R}^{d}$ and used the adaptive choice of the step size and smoothing parameters as proposed in Sec. 3.2. In our experiments, we fixed the number of sections to 50 and 100, with $2048$ and $4096$ bytes each for MalConv and and GBDT-EMBER, resulting in an initial total manipulation amount of $102800$ and $410400$ bytes. It is important to note that additional bytes are needed to include sections inside programs due to format constraints that require content to span multiples of a specific value [12].
@@
Setup of competing attacks. We compare ZEXE different proposed attacks in the literature, by fixing the same amount of queries to 1000.
@@
GAMMA [12]: this algorithm minimizes Eq. 1 with a generic algorithm leveragin the section injection manipulation. GAMMA can not fix a priori the size of the perturbation, but it uses a regularization parameter $\lambda$ to control such a quantity. Thus, we test different values of this parameter, by choosing $\lambda\in[10^{-7},10^{-5},10^{-3}]$. Then we fix the number of sections to inject to $50$ and $100$ when computing the robustness of MalConv and GBDT-EMBER respectively.
@@
OZD [31]: we include in the comparison the zeroth-order algorithm whose ZEXE is based upon. Thanks to our formulation, we can include this algorithm as well, even if it has not been proposed on the domain of Windows malware detection. This method is a structured finite-difference algorithm that uses a gradient surrogate similar to the one in Eq. 7. However, it lacks the explorative procedure of our algorithm. In fact, this algorithm can be considered a special case of Algorithm 1 with $\omega<0$, using the gradient surrogate defined in Eq. 2 and updating the best iterate and function values at any iteration $k$ as in Eq. 6, using the sets $V_{k}:=\{v_{k},v_{k}^{\text{best}}\}$ and $D_{k}:=\{F_{z}(v_{k}),F_{k}^{\text{best}}\}$. For a fair comparison, we used the same sets as ZEXE for the update procedure. Since OZD also leverages finite differences, we set the number of random orthogonal directions used to approximate the gradient to $\ell=5$. We manually tuned its smoothing parameter to $h_{k}=100$ and its step size to $\gamma_{k}=\frac{1}{\sqrt{k}}$ for every iteration $k$. For the encoder, decoder, perturbation, and the injected size, we use the same settings as ZEXE for both models.
@@
Random Search (RS): we include an algorithm that, using section injection, introduces the same amount of sections as the other attacks, with the same perturbation size of ZEXE and OZD. However, the attack randomizes the included content at each query, standing as a baseline for our comparison.
@@
Dataset. To perform the experiments, we built a dataset of $200$ malware divided in different malware classes as shown in Table 1. We retrieve this information by querying VirusTotal, [^4] https:
virustotal.com
and by aggregating the results using AVClass [32] to obtain the most-probable class for each single malware. Among these, we were unable to correctly classify 9 samples among 200.
Evaluation Metrics. For each attack, we report the evasion rate ($ER$), which is the fraction of computed adversarial EXEmples that bypass detection; the mean ($\mu_{\text{queries}}$) and standard deviation ($\sigma_{\text{queries}}$) of the number of queries required to achieve evasion, respectively; the mean ($\mu_{\text{size}}$) and standard deviation ($\sigma_{\text{size}}$) of the number of kilobytes (KB) injected (i.e., the perturbation size); and the mean ($\mu_{\text{conf}}$) and standard deviation ($\sigma_{\text{conf}}$) of the confidence computed on adversarial EXEmples that successfully bypass the target model.
Hardware setup. All experiments have been conducted on workstation equipped with an Intel® Xeon® Gold 5217 CPU with 32 cores running at 3.00 GHz, 192 GB of RAM, and two Nvidia® RTX A6000 GPUs with 48 GB of memory each.

### v2

@@ 3.4 Theoretical Results
In this section, we explain the effect of the exploitation steps on the iteration by providing a theoretical analysis of ZEXE. In particular, we provide convergence rates for non-convex, non-smooth targets. To do so, we consider a simplified setting in which the objective function is assumed to be Lipschitz continuous, i.e., it satisfies the following assumption.
@@
Notice that such an assumption is not satisfied in real-world systems due to the discrete nature of the domain. However, we adopt it here to provide a formal explanation of the iterations of the algorithm and to build intuition about its behavior under controlled conditions. As we observe from Algorithm 1, given a threshold $\omega\geq 0$, for every time-step $k\in\mathbb{N}$, if $\|g_{(G_{k},h_{k})(v_{k})}\|^{2}>\omega$, then the update of the iteration is computed as in structured finite-difference methods. Moreover, as in [32], for $h>0$, $\mathbb{E}_{G}[g_{(G,h)}(\cdot)]=\nabla F_{h}(\cdot)$, where $g_{(G,h)}(\cdot)$ is defined in Eq. 7 and $F_{h}(v):=\mathbb{E}_{u\sim\mathbb{B}^{d}}[F_{z}(v+hu)]$ - see Lemma 2 in Appendix A. Thus, for every $0\leq\underline{K}<\bar{K}$ such that for $k=\underline{K},\ldots,\bar{K}$ the condition $\|g_{(G_{k},h_{k})(v_{k})}\|^{2}>\omega$ holds, we can provide rates for the expectation of the norm of the gradient of the smooth approximation of the objective function $F_{h}$. In the following, we refer to complexity as the number of function evaluations required to achieve an accuracy $\varepsilon\in(0,1)$. We use the following notation,
$\eta_{\underline{K}}^{\bar{K}}=\Bigg(\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h}(x_{i})\|^{2}]\Bigg)/A_{\underline{K}}^{\bar{K}}\quad\text{with}\quad A_{\underline{K}}^{\bar{K}}=\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}.$
In particular, we consider the setting in which the smoothing parameter is chosen as a constant $h_{k}=h$. Now, we state the main theorem for the non-convex non-smooth setting. Proofs are provided in the appendix.
@@
$\eta_{\underline{K}}^{\bar{K}}\leq S_{\underline{K}}^{\bar{K}}/A_{\underline{K}}^{\bar{K}}$
$\text{with}\qquad S_{\underline{K}}^{\bar{K}}=F_{h_{\underline{K}}}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h}.$
@@
Under the same Assumptions of Theorem 3.1, let $\gamma_{k}=\gamma$ with $\gamma,h>0$. Then, for $k=\underline{K},\cdots,\bar{K}$
$\eta_{\underline{K}}^{\bar{K}}\leq\frac{F_{h}(v_{\underline{K}})-\min F_{z}}{\gamma(\bar{K}-\underline{K})}+\frac{L^{3}d^{2}\sqrt{d}}{\ell h}\gamma.$
Moreover, let $K=\bar{K}-\underline{K}$ and $\varepsilon\in(0,1)$. Then, if $K\geq 4(F_{h}(v_{\underline{K}})-\min F_{z})(L^{3}d^{2}\sqrt{d})\varepsilon^{-2}/(\ell h)$, choosing $\gamma=\sqrt{\frac{F_{h}(v_{\underline{K}})-\min F_{z})\ell h}{KL^{3}d^{2}\sqrt{d}}}$, we have that $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$. Thus, the complexity of obtaining $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$ is of the order $\mathcal{O}(d^{2}\sqrt{d}h^{-1}\varepsilon^{-2})$.
Discussion. In Theorem 3.1, we provide a rate on the expected norm of the smoothed gradient. The resulting bound consists of two parts. The first part is due to the (smoothed) functional value at the first point of the sequence for which the norm of the gradient approximation is larger than $\omega$ (i.e., the initialization). The second part is the approximation error. Recall that Assumption 1 holds, and therefore we have that $F_{h}(v)\leq F_{z}(v)+Lh$ for every $v\in\mathbb{R}^{d}$ - see Proposition 1. Thus, by choosing $h$ small, we can reduce the gap between $F_{h}$ and $F$. However, choosing this parameter too small would increase the approximation error significantly. Note that with $\omega=0$, Theorem 3.1 holds for every $k\in\mathbb{N}$. In particular, we obtain a rate that differs from [32, Theorem 2] only in the dependence on the dimension in the second term of $S_{\underline{K}}^{\bar{K}}$. We can achieve the same dependence on the dimension by choosing, for instance, the sequence $h=d\bar{h}$, where $\bar{h}$ is some non-negative sequence. In particular, we observe that if the number of consecutive iterations $k=\underline{K},\ldots,\bar{K}$ for which $\|g_{(G_{k},h_{k})}(v_{k})\|^{2}>\omega$ is sufficiently large, we get that $\eta_{\underline{K}}^{\bar{K}}\leq\varepsilon$. This means that, under such a condition, in expectation, the iterations allow us to approach a stationary point of $F_{h}$.
@@ 4 Experiments
We evaluate the effectiveness of ZEXE by comparing it against GAMMA, a state-of-the-art attack that injects content harvested from benignware samples [12].
Our experimental study is conducted in two distinct settings: comparison with GAMMA and high-dimensional perturbation.
4.1 Comparison with GAMMA
In this first setting, we compare ZEXE with random-search exploration against GAMMA and other approaches using small perturbations.
@@ 4.1 Comparison with GAMMA
Baseline models. We consider two state-of-the-art machine learning Windows malware detectors. The first one is MalConv [31], an end-to-end deep convolutional neural network. MalConv takes in input programs as whole strings of bytes, cropping them if they exceed 2MB of input or, on the contrary, padding them with the special value 256. Due to the absence of any notion of distance in the domain of bytes, these are processed through an 8-dimensional embedding layer that makes them comparable. The second model is GBDT-EMBER [2], which is a gradient-boosting decision tree trained on a complex feature representation introduced to build the open-source EMBER dataset. In particular, EMBER features aggregate information regarding headers, imported APIs, byte histograms, readable strings, and more into 2381 features. For this paper, we leverage an open-source implementation of MalConv, with an input size shrunk to 1MB and detection threshold set to $0.5$, while for GBDT-EMBER we use its original open-source implementation [^3] https://github.com/elastic/ember, with detection threshold set at $0.8$. Both models are wrapped to be attacked through the SecML-Malware Python library [13].
@@
Setup of ZEXE. We set the number of explored directions $\ell=5$ and the threshold on the gradient norm $\omega=10^{-4}$. We set a number of queries of $1000$, thus the number of iterations of ZEXE is $T=1000/(\ell+1)$. As practical manipulation, we use section injection, which includes new content inside input Windows programs, like already-initialized strings, resources, or other assets [12]. We initialize the manipulation with $d$ random printable bytes, ranging from the white-space character (encoded as 32) to the tilde character "~" (encoded as 126). We define the encoding mapping $\psi$ as the function that maps these bytes to their corresponding values in $\mathbb{R}^{d}$. The decoding function $\xi$ projects the perturbed manipulation onto the hypercube $32,126]^{d}\subseteq\mathbb{R}^{d}$. It then takes the integer part of each element in the perturbed manipulation and maps those integers back to the corresponding bytes, ensuring that the bytes remain within the range of printable characters. Finally, the manipulation function $\rho$ modifies the newly injected sections. We defined the sampling space $V$ as the hypercube $32,126]^{d}\subseteq\mathbb{R}^{d}$. The stepsize $\gamma_{k}$ is choosen adaptively as
$\gamma_{k}=\max(\underline{\gamma},\min(F_{z}(v_{k})/\|g_{(G_{k},h_{k})}(v_{k})\|^{2},\bar{\gamma})),$
where $0<\underline{\gamma}\leq\bar{\gamma}$. Moreover, let $\underline{h}>0$ and $h_{0}\geq\underline{h}$. We define the following sequence of smoothing parameters
In our experiments, we fixed the number of sections to 50 and 100, with $2048$ and $4096$ bytes each for MalConv and GBDT-EMBER, resulting in an initial total manipulation amount of $102800$ and $410400$ bytes. It is important to note that additional bytes are needed to include sections inside programs due to format constraints that require content to span multiples of a specific value [12].
@@
Setup of competing attacks. We compare ZEXE different proposed attacks in the literature, by fixing the same number of queries to 1000.
@@
GAMMA [12]: this algorithm minimizes Eq. 1 with a generic algorithm leveraging the section injection manipulation. GAMMA can not fix a priori the size of the perturbation, but it uses a regularization parameter $\lambda$ to control such a quantity. Thus, we test different values of this parameter by choosing $\lambda\in[10^{-7},10^{-5},10^{-3}]$. Then we fix the number of sections to inject to $50$ and $100$ when computing the robustness of MalConv and GBDT-EMBER, respectively.
@@
OZD [32]: we include in the comparison the zeroth-order algorithm whose ZEXE is based upon. Thanks to our formulation, we can include this algorithm as well, even if it has not been proposed in the domain of Windows malware detection. This method is a structured finite-difference algorithm that uses a gradient surrogate similar to the one in Eq. 7. However, it lacks the explorative procedure of our algorithm. In fact, this algorithm can be considered a special case of Algorithm 1 with $\omega<0$, using the gradient surrogate defined in Eq. 2 and updating the best iterate and function values at any iteration $k$ as in Eq. 6, using the sets $V_{k}:=\{v_{k},v_{k}^{\text{best}}\}$ and $D_{k}:=\{F_{z}(v_{k}),F_{k}^{\text{best}}\}$. For a fair comparison, we used the same sets as ZEXE for the update procedure. Since OZD also leverages finite differences, we set the number of random orthogonal directions used to approximate the gradient to $\ell=5$. We manually tuned its smoothing parameter to $h_{k}=100$ and its step size to $\gamma_{k}=\frac{1}{\sqrt{k}}$ for every iteration $k$. For the encoder, decoder, perturbation, and the injected size, we use the same settings as ZEXE for both models.
@@
Random Search (RS): we include an algorithm that, using section injection, introduces the same number of sections as the other attacks, with the same perturbation size of ZEXE and OZD. However, the attack randomizes the included content at each query, standing as a baseline for our comparison. Notice that such a procedure can be recovered as a special case of ZEXE when $\omega\to+\infty$.
4.2 High-dimensional Perturbations
In the second setting, we consider high-dimensional perturbations. More precisely, we focus on comparison ZEXE with different exploration strategies against GAMMA using GBDT-EMBER as detector model. For ZEXE, we set the number of explored directions $\ell=5$. We set a number of queries of $500$. We use the same manipulation of the previous setting, and we initialize the manipulation using $d$ random bytes for each malware, without performing any initialization tuning. We therefore defined the sampling space $V$ as the hypercube $0,256]^{d}\subseteq\mathbb{R}^{d}$. The stepsize $\gamma_{k}$ is chosen with a standard linesearch procedure [4] and the smoothing parameter sequence is selected as in Eq. 10 with $\underline{h}=0.01$, $h_{0}=10^{5}$ and $\omega=10^{-3}$. For every malware in the test set, we set the perturbation size as half of the average perturbation size achieved by GAMMA (with $\lambda=0$) computed over $10$ repetitions. For ZEXE-Langevin (ZEXE-LAN) we set the sequence $\beta_{k}=\frac{1}{2}\frac{\log(k+2)}{k+2}$, while for ZEXE-GA we used a population size of $10$. The threshold on the gradient norm $\omega$ for ZEXE with random search (ZEXE-RS) and ZEXE-GA is set to $10^{-2}$. The number of sections is set to $30$. For GAMMA, we set the regularization parameter $\lambda=0$ and the number of sections to inject to $30,40$, and $50$.
Datasets. To perform the experiments, we built two datasets of $200$ and $700$ malware, divided into different malware classes as shown in Table 1 and Table 2 extracted from the Speakeasy Dataset [38]. We used the first dataset to perform the comparison with state-of-the-art methods and the second to explore the impact of the exploration strategies in high-dimensional setting. We retrieve this information by querying VirusTotal, [^4] https:
virustotal.com
and by aggregating the results using AVClass [35] to obtain the most-probable class for every single malware. Among these, we were unable to correctly classify 9 samples among 200 of the first dataset. For the second dataset, we obtained 700 malware samples by randomly selecting 100 samples from each of the malware families that constitute the Speakeasy dataset.
@@ 4.2 High-dimensional Perturbations
backdoor coinminer dropper keylogger ransomware rat trojan
100 100 100 100 100 100 100
Table 2: Malware dataset used to compare the exploration strategies.
Evaluation Metrics. For each attack, we report the evasion rate ($ER$), which is the fraction of computed adversarial EXEmples that bypass detection; the mean ($\mu_{\text{queries}}$) and standard deviation ($\sigma_{\text{queries}}$) of the number of queries required to achieve evasion, respectively; the mean ($\mu_{\text{size}}$) and standard deviation ($\sigma_{\text{size}}$) of the number of kilobytes (KB) or megabytes (MB) injected (i.e., the perturbation size); and the mean ($\mu_{\text{conf}}$) and standard deviation ($\sigma_{\text{conf}}$) of the confidence computed on adversarial EXEmples that successfully bypass the target model.
Hardware setup. All experiments have been conducted on a workstation equipped with an Intel® Xeon® Gold 5217 CPU with 32 cores running at 3.00 GHz, 192 GB of RAM, and two Nvidia® RTX A6000 GPUs with 48 GB of memory each.